The normals at any point P of the ellipsoid x^2/9 +y^2+4 +z^2 = 1 meet the coordinate planes in Q1,Q2,Q3, respectively. Show that PQ1 : PQ2 : PQ3 :: 9 : 4 : 1.
1
Expert's answer
2020-05-19T18:58:48-0400

The equation of the ellipsoid is "\dfrac{x^2}{9}+\dfrac{y^2}{4}+z^2=1".

Equation of normal at any point "P=(x_0,y_0,z_0)" to the ellipsoid "ax^2+by^2+cz^2=1" is "\frac{x-x_0}{ax_0}=\frac{y-y_0}{by_0}=\frac{z-z_0}{cz_0}".

For the given ellipsoid "a=\frac{1}{9} ; b=\frac{1}{4} ; c=1".

​So, equation of normal at any point "P=(x_0,y_0,z_0)" to the given ellipsoid is

"\frac{9(x-x_0)}{x_0}=\frac{4(y-y_0)}{y_0}=\frac{z-z_0}{z_0}"


When "x=0;" Equation of normal becomes "-9=\frac{4(y-y_0)}{y_0}=\frac{z-z_0}{z_0}"

"\implies y=\frac{-5y_o}{4};z=-8z_o"

Hence, "Q_1=(0,\frac{-5y_o}{4},-8z_o)".

When "y=0; x=\frac{5x_o}{9};z=-3z_o" "\implies Q_2=(\frac{5x_o}{9},0,-3z_o)".

When "z=0;x=8x_o\/9;y=3y_o\/4" "\implies Q_3=(\frac{8x_o}{9},\frac{3y_o}{4},0)".


Hence "PQ_1\n\n\u200b=\\sqrt{(x_o)^2+(y_0-\\frac{-5y_o}{4})^2+(z_0 + 8z_o)^2} = (\\sqrt{16x_o^2+81y_o^2+36^2z_o^2})\/4"

"PQ_2=\\sqrt{(\\frac{4x_o}{9})^2+(y_o)^2+(4z_o)^2} \n\u200b\t\n = (\\sqrt{16x_o^2\u200b +81y_0^2+36^2 z_0^2) }\/9"

and "PQ_3 \u200b=\\sqrt{(x_o - \\frac{8x_0}{9})^2+(y_0-\\frac{3y_o}{4})^2+z_0 ^2} = (\\sqrt{16x_o^2+81y_o^2+36^2z_o^2})\/36" .


Thus "PQ_1\n\n\u200b:PQ_2\n\n\u200b:PQ_3\n\n\u200b=1\/4:1\/9:1\/36"

"\\implies PQ_1\n\n\u200b:PQ_2\n\n\u200b:PQ_3\n\n\u200b=9:4:1"